CH/Agen/97/01

Agenda

Charissa Group Meeting

To be held on: Wednesday, 26 Feb 1997, 11.30 am

To be held at:  Birmingham

Distribution: char-meeting@nnga.daresbury.ac.uk

1. Apologies and welcome
2. ANU    (a) Outstanding work
          (b) When, how and who to carry this out
          (c) Submission to review committee
3. GANIL  (a) Preparations for 12Be and breakup experiments
          (b) DA requirements
          (c) British Council grant
          (d) SPIRAL letter of intent
          (e) EU TMR submission
          (f) Exogam ancillary detectors meeting
          (g) Future proposals
4. STRASBOURG (a) Prep for C+C partial width and 6,7Li fusion experiments
              (b) DA needs
              (c) Future runs and proposals
5. PLANS FOR NEXT 12 MONTHS     (a) Reports from groups
                               (b) Discussion
6. GRANT SPEND                  (a) Discussion
7. PAPERS     (a) Internal circulation of preprints
              (b) Author policy
              (c) MEGHA papers
              (d) Status of other papers
8. ANY OTHER BUSINESS
9. DATE OF NEXT MEETING
      (a) Physics meeting Friday 21 March at Daresbury
      (b) Collaboration meeting (main business to discuss grant spend)

W.N. Catford (Secretary) 5 March 1997